# Omni2: Unifying Omnidirectional Image Generation and Editing in an Omni Model

### v1

@@ Abstract.
$360^{\circ}$ omnidirectional images (ODIs) have gained considerable attention recently, and are widely used in various virtual reality (VR) and augmented reality (AR) applications. However, capturing such images is expensive and requires specialized equipment, making ODI synthesis increasingly important. While common 2D image generation and editing methods are rapidly advancing, these models struggle to deliver satisfactory results when generating or editing ODIs due to the unique format and broad 360 Field-of-View (FoV) of ODIs. To bridge this gap, we construct Any2Omni, the first comprehensive ODI generation-editing dataset comprises 60,000+ training data covering diverse input conditions and up to 9 ODI generation and editing tasks. Built upon Any2Omni, we propose an Omni model for Omni-directional image generation and editing (Omni2), with the capability of handling various ODI generation and editing tasks under diverse input conditions using one model. Extensive experiments demonstrate the superiority and effectiveness of the proposed Omni2 model for both the ODI generation and editing tasks.
@@
copyright: acmlicensedjournalyear: 2025doi: XXXXXXX.XXXXXXXconference: Make sure to enter the correct conference title from your rights confirmation email; October 27–31, 2025; Dublin, Irelandisbn: 978-1-4503-XXXX-X/2018/06
@@
[…]
@@ 1. Introduction
With the rapid advancement of virtual reality (VR) technology, 360 omnidirectional images (ODIs) have gained increasing attention. However, capturing ODIs requires expensive and specialized hardwares, making ODI synthesis a crucial task. While 2D image generation techniques have gradually matured with the rapid advancement of AIGC (Rombach et al., 2022; Chang et al., 2023; Tian et al., 2024), ODI generation remains underexplored. Previous works, such as MVDiffusion (Tang et al., 2023), can only generate ODIs with a limited vertical Field-of-View (FoV) of $360^{\circ}\times 90^{\circ}$, restricting its applicability in real-world scenarios. Though some other methods (Chen et al., 2022; Ye et al., 2024; Zhang et al., 2024) are capable of generating full $360^{\circ}\times 180^{\circ}$ ODIs, they primarily focus on text-driven ODI generation, while ignoring other input conditions such as narrow viewport images, semantic maps, depth maps, etc.
High-quality 2D image editing datasets (Brooks et al., 2023; Zhang et al., 2023; Sheynin et al., 2024) have recently driven the advancement of image editing models (Zhang et al., 2023; Geng et al., 2024). In the context of ODIs, editing is of great importance for enhancing quality of experience in immersive environment. However, omnidirectional image editing remains unexplored. Unlike 2D images, ODIs are stored in the warped equirectangular projection (ERP) format, thus making common 2D image editing algorithms cannot be directly applied. An alternative approach is applying 2D editing models on the designated single view of an ODI after viewport splitting, but it is both time-consuming and inefficient, and may generate inconsistent views. Moreover, due to the unique depth and spatial characteristics of omnidirectional images, conventional 2D editing models struggle to understand the spatial relationships between viewports or even within individual ODI views, making the split-and-edit approach impractical. Therefore, dedicated ODI editing dataset and method are necessary to advance research in this domain.
Built upon the rapid progress in 2D image generation and editing, integrating these tasks into a unified framework has become increasingly popular (Xiao et al., 2024; Xia et al., 2024; Tian et al., 2025). In this paper, we aim to unify ODI generation and editing into an efficient model. To this end, we first introduce Any2Omni, the first comprehensive dataset for omnidirectional image generation and editing tasks. As shown in Table 1, our dataset integrates multiple input modalities for ODI generation tasks. For the newly defined omnidirectional image editing tasks, we start by proposing a simple yet effective pipeline capable of generating high-quality, object-level indoor editing samples. Additionally, we introduce two scene-level ODI editing tasks utilizing existing ODI datasets (Zheng et al., 2020). Overall, our Any2Omni dataset comprises over 60,000 training samples, covering 9 categories of omnidirectional image generation and editing tasks with various input conditions.
Based on Any2Omni, we further introduce the first omni model for omnidirectional image generation and editing, termed Omni2. Omni2 adopts a simple yet effetive Transformer-based framework to support $360^{\circ}\times 180^{\circ}$ high-quality omnidirectional image synthesis under a variety of multimodal input conditions. In contrast to existing diffusion-based ODI generation methods, which incorporate additional attention blocks for multi-view consistency (Tang et al., 2023; Zhang et al., 2024), we introduce a novel approach by executing viewport-based bidirectional attention within a unified Transformer. Our model demonstrates superior performance across a wide range of ODI generation and editing tasks with various input conditions, as shown in Fig. 1.
The main highlights of this work include:
@@ 3.2.1. Object-level Indoor Editing Pipeline
The primary challenge in generating ODI editing datasets lies in creating editing image pairs where only a partial set of objects is precisely added or removed, while the other details of the image are retained. Due to the rarity of ODIs in social media applications, it is not feasible to directly collect editing image pairs through web crawling, as is done with 2D editing datasets (Geng et al., 2024; Sushko et al., 2025). Since constructing manually annotated datasets is impractically time-consuming, 2D image editing datasets are often built by either applying 2D generation methods (Brooks et al., 2023) or utilizing existing 2D-related datasets. For instance, Inst-Inpaint (Yildirim et al., 2023) leverages existing dataset (Hudson and Manning, 2019) to obtain scene graphs within the image, which significantly simplifies the task. However, due to the immaturity of current ODI generation methods and limited sources of ODI-related datasets, we are forced to build the omnidirectional image editing dataset from scratch. In response, we develop a pipeline that can generate high-quality editing image pairs and detailed editing instructions from a single, unconditioned indoor ODI input, as illustrated in Fig. 2.
Compared to outdoor omnidirectional images, indoor ODIs are often more diverse and contain easily segmentable objects. As an initial attempt at omnidirectional image editing, we select Structured3D (Zheng et al., 2020), a large-scale indoor ODI dataset, to generate object-level removal/addition image pairs. Inspired by (Yildirim et al., 2023), we constructed our editing image pairs through a pipeline that integrates segmentation and inpainting models to generate precise edited image pairs. However, unlike Inst-Inpaint, which relies on object categories provided by the scene graphs in the GQA dataset (Hudson and Manning, 2019) for instance segmentation, our task requires an algorithm capable of both detecting instances and outputting corresponding instance categories. To accomplish this, we choose GroundedSAM (Ren et al., 2024) due to its ability to effectively detect and classify instances in a single pass. Directly applying segmentation on the ERP formatted ODIs yields imprecise results. Furthermore, given the richness of objects in indoor scenes, we propose to generate editing pairs with predefined viewports. Specifically, we first split the ODIs into six viewpoints: front, back, left, right, top, and bottom. Then, segmentation is applied on each individual viewport, as shown in Fig. 2(a).
We perform dual-stage post-processing after the segmentation process to ensure data quality, as illustrated in Fig. 2(b). First, we apply instance filtering to remove instances that are irrelevant for our object-level editing purpose, such as “floor”, “cityscape”, and “bedroom”. Since the inpainting operation is applied to each individual viewport, instances that span across multiple viewports may experience disruptions in scene consistency when inpainting is applied to each viewport separately. To address this, we implement an automatic instance filtering method, where instances with segmentation masks located at the image edges are discarded. This process effectively filters out approximately $60\%$ of the segmentation results.
With the selected instance classes and corresponding segmentation masks, we adopt LAMA (Suvorov et al., 2022) to inpaint the instances, thereby generating object-removal outputs, shown in Fig. 2(c)-(d). The outputs are manually selected to ensure image quality. Inspired by (Geng et al., 2024; Yildirim et al., 2023), object-addition pairs are generated by swapping the input and output images. By executing multiple rounds of the pipeline on the same perspective image and stitching together images edited from various viewports, we further obtain a set of multi-object editing image pairs.
Preliminary editing intructions are formed using the detected instance class and the viewport information in the format as “Remove the ¡instance class¿ on my ¡viewport¿.” In certain editing cases, the relative location of the object should be pointed out. To address this, we input both the original image and segmented instance class into Internvl2-5 (Chen et al., 2024), instructing it to generate relative position descriptions, as shown in Fig. 2(e). With these relative position descriptions, we create more detailed editing instructions: “Remove the ¡instance class¿ on my ¡viewport¿ beside the ¡reference instance¿.” For multi-object editing, the instructions are formed as: “Add the ¡instance class1¿ beside the ¡reference instance1¿ on my my ¡viewport1¿ and ¡instance class2¿ on my ¡viewport2¿.” To ensure data diversity, we further utilize Internvl2-5 (Chen et al., 2024) to refine the generated editing instructions, thereby creating a range of diverse expressions that contribute to improving the model’s robustness. The instruction generation process is carried out over multiple iterations, during which we manually select the most appropriate instructions.
@@ 4.2.1. Input Embedding
We propose generating separate viewports instead of treating the entire ODI as a single entity, in order to ensure viewport consistency and align with the capture process of ODIs. To address this, the noise input is defined on a viewport basis for the viewport-based diffusion process, as shown in Fig. 3.
Input image condition is also viewport tokenized to align with the output. Given an ODI as an input image, we employ a viewport tokenizer to convert the image into viewport-based tokens. Specifically, the input ODI is divided into six overlapping viewports, each of which is then transformed into viewport-based latent representations using a frozen VAE encoder. These representations are then flattened into a sequence of visual tokens with the patch size set to 2 following (Peebles and Xie, 2023). During training, the viewport noise is formulated as:
(1) $z^{i}_{t}=tz^{i}+(1-t)\epsilon^{i}$
where $z^{i}$ is the viewport-based latent representation of the ground truth $\{x_{i}\}_{i=1}^{6}$, $t$ is the diffusion timestep and $\epsilon^{i}\sim\mathcal{N}(0,I)$ is the Gaussian noise. During the inference stage, each viewport noise is formulated as a random sampled Gaussian noise, as shown in Fig. 3.
The text prompt and timestep are also tokenized and fed into the pretrained transformer to facilitate the denoising process.
@@ 4.3.2. Loss Function
(2) $\mathcal{L}=\mathbb{E}[\sum_{i=1}^{N}||(z^{i}-\epsilon^{i})-v_{\theta}(z^{i}_{t},t,c)||^{2}]$
where $z^{i}=\{\mathcal{E}(x_{i})\}_{i=1}^{N}$ represents the latent embeddings of the $i$-th viewport image, $\epsilon^{i}\sim\mathcal{N}(0,1)\}_{i=1}^{N}$ is the Gaussian noise, $z^{i}_{t}$ is the noised latent for the $i$-th viewport as defined in Eq. 1, and $c$ denotes the embeded condition.
For tasks like light modify and object-level editing, different regions should be adjusted with varying intensities. Specifically, for light modify, distinct areas require different levels of adjustment, while in object-level editing, only a small portion of the regions needs to be modified, thus we make slight modifications to the loss function for these tasks to enable the model to learn the appropriate intensities and to prevent it from simply copying the input image as the output, as discussed in (Xiao et al., 2024). This was achieved by introducing a weighted loss, where regions that differ significantly from the input image are assigned higher weights. However, although the decoration task belongs to editing, since the change is made to the overall image, applying weighted loss leads to unsatisfatory results. Therefore, we use the original loss for the decoration task.
@@ 5. Experiment
Text2Light (Chen et al., 2022) 91.05 4.90 0.7007 135.36
MVDiffusion* (Tang et al., 2023) 92.59 6.64 0.5367 252.08
PanFusion (Zhang et al., 2024) 80.66 7.36 0.8463 67.83
Omni2 (Ours) 47.32 7.62 0.8887 22.55
5.1. Implementation Details
5.1.1. Training Settings
@@ 5.2.2. Quantitative Results
Table 2 presents the quantitative results of text2ODI task. We use Fréchet Inception Distance (FID) (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016) and Clip Score (CS) (Hessel et al., 2021) to compare the quality of ODIs generated by Omni2 with that by other methods. We also report inference time on a single A6000 for efficiency comparison. It can be observed that our model outperforms others in terms of all the evaluation metrics. Moreover, our method has a substantial advantage in inference time due to the attention mechanism using kv-cache.
@@ 5.2.3. User Study
SIG-SS (Hara et al., 2023) 66.67 5.04 N/A N/A N/A
OmniDreamer (Akimoto et al., 2022) 73.80 5.15 N/A N/A N/A
PanoDiffusion (Wu et al., 2024) 127.30 4.19 N/A N/A N/A
PanoDiff (Wang et al., 2023a) 61.03 6.30 45.54 6.78 0.8535
Omni2 (Ours) 44.13 6.86 37.40 6.93 0.8620
5.3. Multi-modal to Image
The qualitative results of image conditioned outpainting are presented in Fig. 5. We only provide the outpainting results from center viewport here, outpainting results from diverse input masks are presented in the supplementary. It should be noted that since we do not know the specific data split for these models, images in our test set could be in their training set. As can be seen in the figure, outpainting results generated by SIG-SS (Hara et al., 2023) and OmniDreamer (Akimoto et al., 2022) suffer from noticeable boundary inconsistencies, while PanoDiffusion (Wu et al., 2024) tends to produce images with evident artifacts. Although PanoDiff (Wang et al., 2023a) demonstrates relatively better visual quality in terms of boundary consistency, the outpainting content appears less semantically reasonable. In contrast, our model significantly outperforms these baselines by generating semantically meaningful and visually coherent images with superior boundary consistency and enhanced overall visual fidelity.
We also compare the performance of proposed Omni2 with PanoDiff on text-guided ODI outpainting, qualitative results are shown in Fig. 6. PanoDiff struggles to generate semantically plausible results, e.g. , grass in the sky in Fig. 6 (a). In addition, artifacts and meaningless objects appear in the generated ODIs. In contrast, our proposed Omni2 is able to generate meaningful content that is consistent with the text prompt.
Due to the limited number of dedicated methods for ODI inpainting, semantic2image, and depth2image, the qualitative results of these tasks are presented in the supplementary material.

### v2

# Shanghai China

@@ Abstract.
$360^{\circ}$ omnidirectional images (ODIs) have gained considerable attention recently, and are widely used in various virtual reality (VR) and augmented reality (AR) applications. However, capturing such images is expensive and requires specialized equipment, making ODI synthesis increasingly important. While common 2D image generation and editing methods are rapidly advancing, these models struggle to deliver satisfactory results when generating or editing ODIs due to the unique format and broad 360 Field-of-View (FoV) of ODIs. To bridge this gap, we construct Any2Omni, the first comprehensive ODI generation-editing dataset comprises 60,000+ training data covering diverse input conditions and up to 9 ODI generation and editing tasks. Built upon Any2Omni, we propose an Omni model for Omni-directional image generation and editing (Omni2), with the capability of handling various ODI generation and editing tasks under diverse input conditions using one model. Extensive experiments demonstrate the superiority and effectiveness of the proposed Omni2 model for both the ODI generation and editing tasks. Both the Any2Omni dataset and the Omni2 model are publicly available at: https://github.com/IntMeGroup/Omni2.
@@
* Corresponding authors.
copyright: acmlicensedjournalyear: 2025copyright: acmlicensedconference: Proceedings of the 33rd ACM International Conference on Multimedia; October 27–31, 2025; Dublin, Irelandbooktitle: Proceedings of the 33rd ACM International Conference on Multimedia ($M$ ’25), October 27–31, 2025, Dublin, Irelanddoi: 10.1145/3746027.3755405isbn: 979-8-4007-2035-2/2025/10submissionid: 3600ccs: Computing methodologies Computer visionccs: Computing methodologies Virtual reality
@@
Editing Object-level Editing single-object removal/add image+text 6,690 1000
@@ 1. Introduction
With the rapid advancement of virtual reality (VR) technology, 360 omnidirectional images (ODIs) have gained increasing attention. However, capturing ODIs requires expensive and specialized hardwares, making ODI synthesis a crucial task. While 2D image generation techniques have gradually matured with the rapid advancement of AIGC (Rombach et al., 2022; Chang et al., 2023; Liu et al., 2024; Duan et al., 2025; Tian et al., 2024), ODI generation remains underexplored. Previous works, such as MVDiffusion (Tang et al., 2023), can only generate ODIs with a limited vertical Field-of-View (FoV) of $360^{\circ}\times 90^{\circ}$, restricting its applicability in real-world scenarios. Though some other methods (Chen et al., 2022; Ye et al., 2024; Zhang et al., 2024) are capable of generating full $360^{\circ}\times 180^{\circ}$ ODIs, they primarily focus on text-driven ODI generation, while ignoring other input conditions such as narrow viewport images, semantic maps, depth maps, etc. High-quality 2D image editing datasets (Brooks et al., 2023; Zhang et al., 2023; Min et al., 2024; Sheynin et al., 2024; Xu et al., 2025) have recently driven the advancement of image editing models (Zhang et al., 2023; Geng et al., 2024). In the context of ODIs, editing is of great importance for enhancing quality of experience in immersive environment. However, omnidirectional image editing remains unexplored. Unlike 2D images, ODIs are stored in the warped equirectangular projection (ERP) format, thus making common 2D image editing algorithms cannot be directly applied. An alternative approach is applying 2D editing models on the designated single view of an ODI after viewport splitting, but it is both time-consuming and inefficient, and may generate inconsistent views. Moreover, due to the unique depth and spatial characteristics of omnidirectional images, conventional 2D editing models struggle to understand the spatial relationships between viewports or even within individual ODI views, making the split-and-edit approach impractical. Therefore, dedicated ODI editing dataset and method are necessary to advance research in this domain. Built upon the rapid progress in 2D image generation and editing, integrating these tasks into a unified framework has become increasingly popular (Xiao et al., 2024; Xia et al., 2024; Tian et al., 2025). In this paper, we aim to unify ODI generation and editing into an efficient model. To this end, we first introduce Any2Omni, the first comprehensive dataset for omnidirectional image generation and editing tasks. As shown in Table 1, our dataset integrates multiple input modalities for ODI generation tasks. For the newly defined omnidirectional image editing tasks, we start by proposing a simple yet effective pipeline capable of generating high-quality, object-level indoor editing samples. Additionally, we introduce two scene-level ODI editing tasks utilizing existing ODI datasets (Zheng et al., 2020; Yang et al., 2024, 2025; Duan et al., 2022). Overall, our Any2Omni dataset comprises over 60,000 training samples, covering 9 categories of omnidirectional image generation and editing tasks with various input conditions. Based on Any2Omni, we further introduce the first omni model for omnidirectional image generation and editing, termed Omni2. Omni2 adopts a simple yet effetive Transformer-based framework to support $360^{\circ}\times 180^{\circ}$ high-quality omnidirectional image synthesis under a variety of multimodal input conditions. In contrast to existing diffusion-based ODI generation methods, which incorporate additional attention blocks for multi-view consistency (Tang et al., 2023; Zhang et al., 2024), we introduce a novel approach by executing viewport-based bidirectional attention within a unified Transformer. Our model demonstrates superior performance across a wide range of ODI generation and editing tasks with various input conditions, as shown in Fig. 1. The main highlights of this work include:
@@ 3.2.1. Object-level Indoor Editing Pipeline
The primary challenge in generating ODI editing datasets lies in creating editing image pairs where only a partial set of objects is precisely added or removed, while the other details of the image are retained. Due to the rarity of ODIs in social media applications, it is not feasible to directly collect editing image pairs through web crawling, as is done with 2D editing datasets (Geng et al., 2024; Sushko et al., 2025). Since constructing manually annotated datasets is impractically time-consuming, 2D image editing datasets are often built by either applying 2D generation methods (Brooks et al., 2023) or utilizing existing 2D-related datasets. For instance, Inst-Inpaint (Yildirim et al., 2023) leverages existing dataset (Hudson and Manning, 2019) to obtain scene graphs within the image, which significantly simplifies the task. However, due to the immaturity of current ODI generation methods and limited sources of ODI-related datasets, we are forced to build the omnidirectional image editing dataset from scratch. In response, we develop a pipeline that can generate high-quality editing image pairs and detailed editing instructions from a single, unconditioned indoor ODI input, as illustrated in Fig. 2. Compared to outdoor omnidirectional images, indoor ODIs are often more diverse and contain easily segmentable objects. As an initial attempt at omnidirectional image editing, we select Structured3D (Zheng et al., 2020), a large-scale indoor ODI dataset, to generate object-level removal/addition image pairs. Inspired by (Yildirim et al., 2023), we constructed our editing image pairs through a pipeline that integrates segmentation and inpainting models to generate precise edited image pairs. However, unlike Inst-Inpaint, which relies on object categories provided by the scene graphs in the GQA dataset (Hudson and Manning, 2019) for instance segmentation, our task requires an algorithm capable of both detecting instances and outputting corresponding instance categories. To accomplish this, we choose GroundedSAM (Ren et al., 2024) due to its ability to effectively detect and classify instances in a single pass. Directly applying segmentation on the ERP formatted ODIs yields imprecise results. Furthermore, given the richness of objects in indoor scenes, we propose to generate editing pairs with predefined viewports. Specifically, we first split the ODIs into six viewpoints: front, back, left, right, top, and bottom. Then, segmentation is applied on each individual viewport, as shown in Fig. 2(a). We perform dual-stage post-processing after the segmentation process to ensure data quality, as illustrated in Fig. 2(b). First, we apply instance filtering to remove instances that are irrelevant for our object-level editing purpose, such as “floor”, “cityscape”, and “bedroom”. Since the inpainting operation is applied to each individual viewport, instances that span across multiple viewports may experience disruptions in scene consistency when inpainting is applied to each viewport separately. To address this, we implement an automatic instance filtering method, where instances with segmentation masks located at the image edges are discarded. This process effectively filters out approximately $60\%$ of the segmentation results. With the selected instance classes and corresponding segmentation masks, we adopt LAMA (Suvorov et al., 2022) to inpaint the instances, thereby generating object-removal outputs, shown in Fig. 2(c)-(d). The outputs are manually selected to ensure image quality. Inspired by (Geng et al., 2024; Yildirim et al., 2023), object-addition pairs are generated by swapping the input and output images. By executing multiple rounds of the pipeline on the same perspective image and stitching together images edited from various viewports, we further obtain a set of multi-object editing image pairs. Preliminary editing intructions are formed using the detected instance class and the viewport information in the format as “Remove the ¡instance class¿ on my ¡viewport¿.” In certain editing cases, the relative location of the object should be pointed out. To address this, we input both the original image and segmented instance class into Internvl2-5 (Chen et al., 2024), instructing it to generate relative position descriptions, as shown in Fig. 2(e). With these relative position descriptions, we create more detailed editing instructions: “Remove the ¡instance class¿ on my ¡viewport¿ beside the ¡reference instance¿.” For multi-object editing, the instructions are formed as: “Add the ¡instance class1¿ beside the ¡reference instance1¿ on my my ¡viewport1¿ and ¡instance class2¿ on my ¡viewport2¿.” To ensure data diversity, we further utilize Internvl2-5 (Chen et al., 2024) to refine the generated editing instructions, thereby creating a range of diverse expressions that contribute to improving the model’s robustness. The instruction generation process is carried out over multiple iterations, during which we manually select the most appropriate instructions.
@@ 4.2.1. Input Embedding
We propose generating separate viewports instead of treating the entire ODI as a single entity, in order to ensure viewport consistency and align with the capture process of ODIs. To address this, the noise input is defined on a viewport basis for the viewport-based diffusion process, as shown in Fig. 3. Input image condition is also viewport tokenized to align with the output. Given an ODI as an input image, we employ a viewport tokenizer to convert the image into viewport-based tokens. Specifically, the input ODI is divided into six overlapping viewports, each of which is then transformed into viewport-based latent representations using a frozen VAE encoder. These representations are then flattened into a sequence of visual tokens with the patch size set to 2 following (Peebles and Xie, 2023). During training, the viewport noise is formulated as:
(1) $\vskip-3.0ptz^{i}_{t}=tz^{i}+(1-t)\epsilon^{i}\vskip-3.0pt$
where $z^{i}$ is the viewport-based latent representation of the ground truth $\{x_{i}\}_{i=1}^{6}$, $t$ is the diffusion timestep and $\epsilon^{i}\sim\mathcal{N}(0,I)$ is the Gaussian noise. During the inference stage, each viewport noise is formulated as a random sampled Gaussian noise, as shown in Fig. 3. The text prompt and timestep are also tokenized and fed into the pretrained transformer to facilitate the denoising process.
@@ 4.3.2. Loss Function
[…]
where $z^{i}=\{\mathcal{E}(x_{i})\}_{i=1}^{N}$ represents the latent embeddings of the $i$-th viewport image, $\epsilon^{i}\sim\mathcal{N}(0,1)\}_{i=1}^{N}$ is the Gaussian noise, $z^{i}_{t}$ is the noised latent for the $i$-th viewport as defined in Eq. 1, and $c$ denotes the embeded condition. For tasks like light modify and object-level editing, different regions should be adjusted with varying intensities. Specifically, for light modify, distinct areas require different levels of adjustment, while in object-level editing, only a small portion of the regions needs to be modified, thus we make slight modifications to the loss function for these tasks to enable the model to learn the appropriate intensities and to prevent it from simply copying the input image as the output, as discussed in (Xiao et al., 2024). This was achieved by introducing a weighted loss, where regions that differ significantly from the input image are assigned higher weights. However, although the decoration task belongs to editing, since the change is made to the overall image, applying weighted loss leads to unsatisfatory results. Therefore, we use the original loss for the decoration task.
@@ 5. Experiment
Text2Light (Chen et al., 2022) 2.70 91.05 4.90 0.7007 135.36
MVDiffusion* (Tang et al., 2023) 3.06 92.59 6.64 0.5367 252.08
PanFusion (Zhang et al., 2024) 2.54 80.66 7.36 0.8463 67.83
SD+LoRA (Rombach et al., 2022) 2.30 57.97 7.41 0.8540 31.25
Omni2 (Ours) 2.25 47.32 7.62 0.8887 22.55
@@ 5.2.2. Quantitative Results
Table 2 presents the quantitative results of text2ODI task. We also LoRA-finetuned Stable Diffusion (Rombach et al., 2022) to provide more comparative data. We use Fréchet Auto-Encoder Distance (FAED) (Oh et al., 2022), Fréchet Inception Distance (FID) (Heusel et al., 2017), Inception Score (IS) (Salimans et al., 2016) and Clip Score (CS) (Hessel et al., 2021) to compare the quality of ODIs generated by Omni2 with that by other methods. We also report inference time on a single A6000 for efficiency comparison. It can be observed that our model outperforms others in terms of all the evaluation metrics. Moreover, our method has a substantial advantage in inference time due to the attention mechanism using kv-cache.
@@ 5.3. Multi-modal to Image
SIG-SS (Hara et al., 2023) 1.08 66.67 5.04 N/A N/A N/A N/A
OmniDreamer (Akimoto et al., 2022) 1.83 73.80 5.15 N/A N/A N/A N/A
PanoDiffusion (Wu et al., 2024) 1.54 127.30 4.19 N/A N/A N/A N/A
PanoDiff (Wang et al., 2023a) 2.21 61.03 6.30 1.22 45.54 6.78 0.8535
Omni2 (Ours) 1.00 44.13 6.86 0.99 37.40 6.93 0.8620
The qualitative results of image conditioned outpainting are presented in Fig. 5. We only provide the outpainting results from center viewport here, outpainting results from diverse input masks are presented in the supplementary. It should be noted that since we do not know the specific data split for these models, images in our test set could be in their training set. As can be seen in the figure, outpainting results generated by SIG-SS (Hara et al., 2023) and OmniDreamer (Akimoto et al., 2022) suffer from noticeable boundary inconsistencies, while PanoDiffusion (Wu et al., 2024) tends to produce images with evident artifacts. Although PanoDiff (Wang et al., 2023a) demonstrates relatively better visual quality in terms of boundary consistency, the outpainting content appears less semantically reasonable. In contrast, our model significantly outperforms these baselines by generating semantically meaningful and visually coherent images with superior boundary consistency and enhanced overall visual fidelity. We also compare the performance of proposed Omni2 with PanoDiff on text-guided ODI outpainting, qualitative results are shown in Fig. 6. PanoDiff struggles to generate semantically plausible results, e.g. , grass in the sky in Fig. 6 (a). In addition, artifacts and meaningless objects appear in the generated ODIs. In contrast, our proposed Omni2 is able to generate meaningful content that is consistent with the text prompt. Due to the limited number of dedicated methods for ODI inpainting, semantic2image, and depth2image, the qualitative results of these tasks are presented in the supplementary material.